Linux Commercial-HOWTO

Harald Milz, hm@ix.de TODAY This document contains commercial software and applications which are offered for Linux. It is maintained by

Harald Milz (http://www.ix.de/editors/hm.html) (hm@ix.de).

Contents

1 Introduction 2 2 Disclaimer 4 3 Related Information 4 4 Marketing aspects of commercial Linux applications 4 5 Product Groups 4

5.1 Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.1.1 D-ISAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.1.2 Flagship . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.1.3 Just Logic/SQL Database Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 5.1.4 POET 2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.1.5 Veritas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.6 Yard SQL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5.2 Data Visualisation, CAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.2.1 Ghostscript 3.x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.2.2 MRJ Symbolic OCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.2.3 SISCAD-P 1.3-3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.2.4 TecPlot 6.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

5.3 Development tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.3.1 Basmark QuickBASIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.3.2 ISE Eiffel 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.3.3 Metacard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.3.4 tgdb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.4 Text Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

5.4.1 CRISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.4.2 ibgsXaed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

5.5 X Windows related stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.5.1 Metrolink Motif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.5.2 X Inside X Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

5.6 Other Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.6.1 aqua zis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.6.2 Maple V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.6.3 Reduce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 5.6.4 ROSIN NC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

5.7 Free Software for Commercial Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.7.1 Cyclades Cyclom-Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

1 Introduction

This is the Linux Commercial-HOWTO. It contains commercial application software which is available for Linux. The Commercial-HOWTO will not contain Linux distributions and related stuff – that’s what the Distribution-HOWTO is for. This HOWTO will contain tabular entries for each product. The entry format is similar to the new Linux Software Map (LSM) entry (field/stanza lengths are arbitrary). The sample entries are just examples. Please: KEEP SHORT. Otherwise I’ll have to shorten your data, and I hate work. ;-)

Description:

SHORT description of the package, just the basic functionality Distribution Media:

diskettes, CD-ROM, tapes or others Licensing Policy:

whatever applies. is there a free demo or shareware version available via FTP? Where? OS Provisions:

kernel version, XFree86 version, Motif version, RAM and harddisk usage Documentation:

printed documentation, page number, online help, language Product Support:

update service, maintenance, training and prices thereof

Linux Support:

if you bundle a complete Linux system with your software, which kind of support do you offer for the operating system? prices?

Extra Features and Add-Ons:

and their prices

Available Since:

the date you started shipping the Linux version

Countries with Distribution:

if you ship in too many countries to name them all, naming regional distribution channels is okay.

Price Range:

whatever it costs

Installed Base:

how many copies of the Linux version did you sell?

Vendor:

          address
          phone
          fax
          e-mail address
          contact person

I’ll post the list to several national and international newsgroups on a monthly basis. In addition, the Commercial-HOWTO will be available via FTP and FSP (port 21): ftp.ix.de:/pub/Linux/docs/HOWTO (the latest and greatest) and uploaded to sunsite.unc.edu:/pub/Linux/docs/HOWTO weekly if there were changes. WWW users

can access the document at URL ftp://ftp.ix.de/pub/Linux/docs/HOWTO/Commercial-HOWTO.html.

There is a french version of sections 1, 2 and 3 available via WWW

(ftp://ftp.ix.de/pub/Linux/docs/HOWTO/Commercial-HOWTO.French.html).

The Commercial-HOWTO is not a forum for product announcements or marketing hypes but a service to customers primarily. Two main goals are being aimed at:

1. It shall help companies who want to run Linux to find software solutions and applications. The international distribution of this list will enhance the contact opportunity.

2. It is meant to prove the commercial useability of Linux and thus to encourage other vendors to port their software as well.

Companies and developers who are offering their products for Linux and interested in joining the Commercial-HOWTO are invited to fill in the following form and contact me via e-mail (hm@ix.de, preferred) or fax (+49 (0)511/5352-378). If you happen to know a company which does, please drop me a note. If you send me an entry for your product(s), please don’t forget to send me timely updates if appropriate.

2 Disclaimer

I will not select nor deselect any particular product. Instead, everyone who wants to have his/her product included will be serviced. However, I reserve the right to shorten individual entries to keep things concise.

If you don’t find a particular product or vendor in this list, this is probably due to one of the following reasons:

• I never heard of that product or vendor and thus didn’t try to get in contact.

• I did get in contact, but the vendor didn’t answer yet.

• I did get in contact, but the vendor stated positively that he doesn’t sell his product for Linux (yet).

In any case, please kick me if you feel someone’s missing, also if you discover any errors in the file. But don’t forget: I’m not being paid for that job ;-) .

Sometimes two vendor’s addresses are mentioned in the ”Vendor:” field. In these cases I received the information from the German subsidiary / distributor. The original manufacturer’s address is always mentioned first.

3 Related Information

There’s another html page which covers commercial Linux software too. It is provided by LINUX.ORG.UK

(http://www.linux.org.uk/LxCommercial.html) and maintained by Alan Cox (Alan.Cox@linux.org).

4 Marketing aspects of commercial Linux applications

This section was suggested by Orest Zborowski (orestz@eskimo.com) and is still under construction. It is meant to help potential commercial software developers to market their products. There are several caveats to watch when doing this, e.g. the GNU Public License (ftp://ftp.ix.de/pub/Linux/COPYING-2.0). I’d appreciate anyone who wants to share his/her experience with Linux-specific marketing issues to send me some lines. Please make sure that people interested in more specific items can contact you via e-mail or fax.

5 Product Groups

This list is divided into several product groups to make finding a specific product more easy.

5.1 Databases

Database products are well appreciated, even under Linux ;-)

5.1.1 D-ISAM Description:

Multikey B+ tree Isam File Handler. Follows the C-ISAM (Informix) file structure and function calls. Sold with Source.

Distribution Media:

5.25” and 3.5” Floppy disk Licensing Policy:

Free distribution of the first 100 D-ISAM’d executables. Royalties apply after that. OS Provisions:

Independent Documentation:

Documentation in english. Unix Style manual. Product Support:

email, bbs, fax, telephone. First 90 days free. US$450/year for ongoing. Linux Support:

N/A Extra Features and Add-Ons:

Transaction Processing functions: US$295.00 Available Since:

June 1994 Countries with Distribution:

UK, Germany, USA, Canada (manufacturer) Price Range:

Single Programmer License with Source: US$595, Transaction Processing Add-on: US$295, Additional Programmers: US$100, Runtime Licenses per 100: US$695 (after the first 100), Unlimited Runtime: US$6950. A free D-ISAM for Linux library is available for development, but not for commercial use,

     via ftp (ftp://ftp.wimsey.com/pub/linux/d-isam)

Installed Base: Vendor:

     Byte Designs Ltd
     20568 - 32 Avenue
     Langley, BC V3A 4P5, Canada
     Phone:  +1 (604) 534-0722
     Fax:    +1 (604) 534-2601
     Contact: Heinz Wittenbecher
     E-mail: sales@byted.com

Entered:

Sep 08, 1994

5.1.2 Flagship

Description:

CA-Clipper5, Fox, dBase & beyond for Unix. XBase 4GL applications development system & database. Superset of CA-Clipper. Can be used to port.

Distribution Media:

3 3.5” Diskettes, other media extra charge. Licensing Policy:

No royalty or licensing for software you create and distribute. For more information available via ftp

(ftp://ftp.wgs.com/pub2/wgs/Filelist)

OS Provisions:

Linux 0.99 and later, has been tested on many. Requires C compiler, libraries and linker. Make and other utilities useful. 4 MByte RAM & 5 MByte Hard disk needed to run. More is better.

Documentation:

1200 page printed manual extra charge, same manual online is shipped as part of the product. Product Support:

Support via phone, email, fax (email preferred), unlimited for 30 days, chargable thereafter. Maintenance agreements including discounts on upgrades are available. All of this is country dependent.

Linux Support:

Country Dependent Extra Features and Add-Ons:

FS2 Toolbox, FoxKit, and others. Country dependent. Available Since:

July 1994 Countries with Distribution:

We sell to all countries, and have regional distributors in most major areas, and on all continents.

Price Range:

Country dependent. North America: Single user US$199, Unlimited User US$499 until Nov 15, 1994. US$499 Single & US$999 Unlimited after that.

Installed Base:

Vendor:

     multisoft Datentechnik GmbH
     PO Box 312
     82027 Gruenwald, Germany
     Phone:  +49 (0)89/6417904
     Fax:    +49 (0)89/6412974
     E-Mail: 100031.267@compuserve.com
     Contact:  Dorte Balek
     North America:
     WorkGroup Solutions, Inc
     PO Box 460190
     Aurora, CO 80046-0190, USA
     Phone:  +1 (303) 699-7470
     Fax:    +1 (303) 699-2793
     E-Mail: info@wgs.com
     Contact: Virginia Lane

Entered:

Sep 19, 1994

5.1.3 Just Logic/SQL Database Manager

Description:

The Just Logic/SQL Database Manager is a complete ANSI SQL database system. It includes a complete set of libraries, utilities and database engine for developing C and C++ application. It comes with a choice of three programming interfaces: a C precompiler, a C API interface and a C++ SQL class. In november Just Logic Technologies will release a client-server version permitting Windows, Mac and Unix applications to access a remote server on Linux across TCP.

Distribution Media:

Diskettes.

Licensing Policy:

Unlimited runtime licence of applications.

OS Provisions:

SLS 1.0, 300 K RAM and 1.1 Meg Hard disk space usage.

Documentation:

User reference manual (150 pages) + Unix version reference guide. Product Support:

Phone hotline. ($2.00 per minute charge for technical support, first minute is free.) Linux Support:

same. Extra Features and Add-Ons:

Fully compatible versions on other UNIXes, DOS, Windows and OS/2. Available Since:

August 94 Countries with Distribution:

The product is shipped from the Montreal office to all countries. Price Range:

The Linux release is priced at 195.00 $US plus $10 shipping and handling to US and Canada ($35 other countries)

Installed Base: Vendor:

          Just Logic Technologies Inc.
          P.O. Box 63050
          40 Commerce St.
          Nun's Island, Quebec H3E 1V6, Canada
          Phone:  +1 (800) 267-6887 (toll free USA and Canada)
                  +1 (514) 761-6887
          Fax:    +1 (514) 642-6480 (has voice-fax autodetection)
          E-mail: 71563.3370@CompuServe.COM

Entered:

Sep 08, 1994

5.1.4 POET 2.1

Description:

POET 2.1 ODBMS for C++. Database functionality for C++ objects. Full support of encapsulation, inheritance, polymorphism and objectidentity. Two versions available : Personal Edition (Single/User 1 Developer) Professional Edition (Client/Server up to 4 Developer)

Distribution Media:

Diskettes Licensing Policy:

SDK license from US$ 99. Runtime Licenses for commercial Products and for Public Domain or Share Products available.

OS Provisions:

kernel version > 1.0; SDK needs 4 MBytes on harddisk Documentation:

handbook and tutorial (450+ pages) Product Support:

hotline support, training please contact vendor Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

March 1994 Countries with Distribution:

many; please contact vendor. Price Range:

Personal Edition US$ 99 / DM 170. Professional Edition US$ 1990 / DM 3.990 (includes 1 year hotline support)

Installed Base:

234 copies as of July 1994 Vendor:

          POET Software GmbH
          Fossredder 12
          D-22359 Hamburg, Germany
          Phone:  +49 (0)40/609 90 18
          Fax:    +49 (0)40/603 98 51
          E-Mail: info@poet.de
          Contact: Detlef Meyer

Entered:

Sep 19, 1994 5.1.5 Veritas

Description:

Veritas is a system for German drinks wholesalers (”Getränkegrosshändler”). It’s based on Onyx which is my 4gl and database project, which of course stays under the GNU Public License 2.0 (ftp://ftp.ix.de/pub/Linux/COPYING-2.0). Ftp to wowbagger.pc-labor.uni-bremen.de

(ftp://wowbagger.pc-labor.uni-bremen.de/pub/unix/databases) if you like to take a look at

Onyx. The first version of Veritas was developed for PCOS Olivetti in 1992, later Xenix versions followed. The recent 3rd release features twelve years of experience with this kind of market, so you can expect to have nearly anything a drinks wholesaler needs, e.g. Pfand, Sektsteuer, Brauereiabrechung, Bruchvergütung, etc. Several servers can be linked via SLIP or UUCP and replicate transactions to share data. Clients are connected via ethernet to a local server and can run MS Windows or better Linux.

Distribution Media:

Directly installed on the customer’s hardware. Plus a Tape containing (I hope) all of the GNU’ish sources including Onyx and a bunch of different databases.

Licensing Policy:

Veritas is licensed according to a German translation of the GNU Public License 2.0 (http://www.suse.de/gpl-ger.html), restricted in a way that any changes made must be mailed to the author, and that any customer must know that I’m the author and give support. Although the licence is very much like the GPL, I really recommend to purchase a copy if you prefer to get support.

OS Provisions:

Linux 1.0.9 (MCC), XFree86 2.1. Servers: 8-16 MB RAM, 200-500 MB harddisk, tape, modem. Clients: 8 MB RAM, 40 MB harddisk.

Documentation:

About 41 pages printed for Veritas in German, and about 43 pages online documentation for Onyx in English.

Product Support:

Retail price includes one week of training, and one year first class support over modem lines and D2 phone directly from the author. If you are interested in Onyx support, contact the author directly.

Linux Support:

Support for the installed sites is included in the installation price for one year.

Extra Features and Add-Ons:

N/A Available Since:

The 3rd version of veritas is now in BETA in Bremerhaven. Countries with Distribution:

Only distributed in Germany because of the user interface language and a lot of German tax related features. If sold outside the northern part of Germany we need a VAR able to handle Linux, and sell and install the right hardware for a share.

Price Range:

Between 10,000 and 20,000 DM without hardware which depends on the number of servers and additional tayloring.

Installed Base: Vendor:

          Michael Koehne
          Loss Datensysteme
          Bremerstr. 117
          D-28816 Brinkum, Germany
          Phone:  +49 (0)421/87 55 00
          V32bis: +49 (0)421/87 05 32
          Fax:    +49 (0)421/87 55 51
          E-Mail: kraehe@nordwest.de (company)
                  kraehe@bakunin.north.de (private)

Entered:

Sep 08, 1994

5.1.6 Yard SQL

Description:

The YARD company offers 5 SQL products as follows:

• YARD-SQL – Relational SQL database server with compliance to X/Open XPG4 and ANSI SQL 92

• YARD-ESQLC – Embedded SQL for C.

• YARD-ODBC – ODBC interface for MS Windows clients

• YARD-NET – Remote access to YARD databases via TCP/IP.

• YARD-X – Motif client for database access (No development tool) Distribution Media:

Floppy disk and streamer tapes Licensing Policy:

License number and activation key for each product and installation with user dependent licenses for the SQL-Server.

OS Provisions:

LINUX 1.0.9 and upwards compatible versions; Motif 1.2.3; RAM Usage: 1 MByte (minimum) for Shared Memory, 500 KByte per User; Disk Usage: YARD-SQL 7 MByte, YARD-ESQLC 1 MByte, YARD-ODBC 1,5 MByte, YARD-NET 1 MByte, YARD-X 5 MByte

Documentation:

Printed 600 pages reference and users guide in German. English documentation will be available early 1995.

Product Support:

Release Update, Price 25% Version Update, Price 50% Technical support according to a special maintenance agreement (20% of total pricing). Training on request

Linux Support:

none Extra Features and Add-Ons:

All products also available for other widely distributed Unix systems (e.g. SCO UNIX, SPARC Solaris). Available Since:

January 1994 Countries with Distribution:

Current distribution in Germany through Distributor Ordix Software (info@ordix.de), Paderborn, Germany. Europeean/international distribution scheduled for early 1995.

Price Range:

YARD-SQL 590,00 (1 user) - 8.490,00 DEM (unlimited user) YARD-ESQLC 390,00 DEM YARD-ODBC 490,00 DEM YARD-NET 490,00 DEM YARD-X 990,00 DEM

Installed Base: Vendor:

     YARD Software GmbH
     Hansestr. 99
     51149 Koeln, Germany
     Phone:  +49 (0)22 03/45 71 30
     Fax:    +49 (0)22 03/45 71 31
     E-Mail: yard@yard.de
     Contact: Thomas Schonhoven (thomass@yard.de)

Entered:

Sep 24, 1994

5.2 Data Visualisation, CAD

CAD tools, renderers, OCR software and such stuff.

5.2.1 Ghostscript 3.x

Description:

Aladdin Ghostscript 3.n is a full PostScript Level 2 language interpreter. It can display PostScript files on the screen with X Windows, convert them to various raster formats (TIFF/F, GIF, PCX, PPM), and print them on many non-PostScript printers, such as the H-P inkjet and laser printers.

Distribution Media:

PC diskettes only, normally compressed with tar + gzip. A shareware version is available via ftp

(ftp://ftp.cs.wisc.edu/pub/ghost/aladdin) for non-commercial use only.

Licensing Policy:

Redistribution in non-commercial contexts (i.e., if no money changes hands in connection with the distribution) is allowed without license. Any commercial distribution of any kind requires a commercial license, with a negotiated license fee. If you are a commercial organization and intend to distribute Aladdin Ghostscript in a manner for which you believe you don’t need a license, please check with us: ”in connection with” is only a shorthand for the terms of the actual non-commercial license.

OS Provisions:

Will run on any Linux version. Will display on any X11R4 or R5 compatible window system. Documentation:

English, probably about 30 pages. On-line only (no printed doc). No on-line help. Product Support:

No training available. Maintenance (bug fixing) and updates are available with commercial licenses and unavailable otherwise.

Linux Support:

N/A Extra Features and Add-Ons:

We offer commercial-quality sets of the 35 standard PostScript fonts at a price of $10 to end users and a much lower price to commercial licensees.

Available Since:

August 1994. Countries with Distribution:

We distribute worldwide. Our only current overseas distributor is in Germany.

Price Range:

We will ship diskettes to end-users for a charge of $58 to cover service and materials costs. There is an additional charge for $50 for a custom configuration and of $20 (domestic) or $30 (international) for next-day mail. Commercial licensing prices are negotiated, and vary considerably.

Installed Base:

(L. Peter Deutsch estimates the number of users of the commercial version under Linux at about 12,000, as of Sep 23, 1994.)

Vendor:

     Aladdin Enterprises
     203 Santa Margarita Ave.
     Menlo Park, CA 94025, USA
     Phone:  +1 (415) 322-0103
     Fax:    +1 (415) 322-1734
     E-mail: ghost@aladdin.com

Entered:

Aug 26, 1994

5.2.2 MRJ Symbolic OCR

Description:

MRJ SOCR (Symbolic OCR) (TM) is an OCR program for Japanese text. SOCR reads TIFF files and recognizes Japanese text in scanned images. Output formats include PC, Mac, and Unix formats with Unicode, Shift-JIS, JIS, and EUC encodings. Also available for Sun OS.

Distribution Media:

3.5” Floppy Disks. Licensing Policy:

One SOCR license per workstation. OS Provisions:

Operating System: Linux 1.x, Window System: Motif run time library required, System RAM: 16 MB, Hard Disk: 6 MB.

Documentation:

Installation Guide and User Manual. All documentation is in English. Product Support:

Technical support available via telephone, e-mail, and FAX. Linux Support:

The Linux version of SOCR is fully supported. See Product Support.

Extra Features and Add-Ons:

Contact MRJ for information on OEM and software developer versions of SOCR.

Available Since:

August 1994

Countries with Distribution:

Distribution from U.S. Available for export.

Price Range:

Special Introductory Offer – $895 through 30 September 1994.

Installed Base:

Vendor:

     MRJ, Inc.
     10455 White Granite Drive
     Oakton, VA 22124, USA
     Phone:  +1 (703) 385-0700
     Fax:    +1 (703) 385-4637
     E-Mail: socr@mrj.com

Entered:

Sep 08,1994

5.2.3 SISCAD-P 1.3-3

Description:

SISCAD-P is a 2D-CAD system that gives engineers a production implementation of new design technologies - parametrics, variational geometry, inference sketching, a fully customizable user interface, constraint-based modeling and feature-based modeling. Today there is only a German version of SISCAD-P available. Unfortunately we can not yet provide a SISCAD-P version in any foreign language but we are already working on an English version. The release date for this version is not certain yet.

Distribution Media:

available  via  ftp

(sunsite.unc.edu://pub/Linux/X11/xapps/graphics/draw/siscadp1.3-3.tar.gz). On diskette

and CD-ROM available from several Linux distributors (i.e. CDs with sunsite mirrors. hm)

Licensing Policy:

SISCAD-P is running on several Unix workstations and is available as shareware for Linux. The shareware version is a FULL featured package which is restricted to at maximum 10 parts with each 2000 objects. It is provided at no charge to the user for evaluation. With this concept, the user can test all features of SISCAD-P and has a chance to try before buying it. After a reasonable trial period, the user must make a registration payment and will receive a license code which will disable all limitations.

OS Provisions:

X Windows running under Linux. 8 MByte of memory minimum, 16 MByte or more recommended. Slackware Linux 0.19pl14 or newer. Fvwm or Mwm window manager recommended.

Documentation:

Online help and several documents in PostScript, DVI and ASCII format. Product Support:

Product support for registered users via e-mail, fax and BBS Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

February 1994 Countries with Distribution:

Germany (worldwide via ftp). Price Range:

DM 399 commercial users; DM 49 students version. Installed Base:

20 registered users as per Sep 22, 1994. Vendor:

          Staedtler Mars GmbH & Co
          Geschaeftsbereich Informationssysteme - SIS
          Moosaeckerstrasse 3
          D-90427 Nuernberg, Germany
          Phone: +49 (0)911/3080-691
          Fax:   +49 (0)911/3080-692
          BBS:   +49 (0)911/3080-609, login: info, password: gast
          E-Mail: support@SIS.Staedtler.DE
          Contact: Helmrich Streitmatter

Entered:

Sep 19, 1994 5.2.4 TecPlot 6.0 Description:

Tecplot is a plotting program for visualizing and analyzing engineering and scientific data. The standard version includes XY, 2D and 3D-surface plotting. An optional extension (called 3DV) adds the capability to visualize 3D volumetric data. With Tecplot you don’t have to write a program, just input your data and start visualizing immediately. (for more information consult vendor. hm)

Distribution Media:

3.5-inch floppies, 150MB-QIC, 4mm DAT, 8mm tape, (CD-ROM available November 1994) Licensing Policy:

Single-computer license (node-locked) Evaluation copies are available. OS Provisions:

4 MByte is minimum RAM, required hard disk space is 20 MByte. Documentation:

Printed Users Manual 512 pages, partial online help info. English. Product Support:

Update and technical support included with license for first three months and available for extra charge after that. Training classes are available; price is negotiable.

Linux Support:

N/A Extra Features and Add-Ons:

3DV is an optional package of volumetric features available for an additional $300 USD. Available Since:

August 1993 Countries with Distribution:

(many, please consult vendor. hm) Price Range:

$995 to $1395 USD per Node-locked License Installed Base: Vendor:

          Amtec Engineering, Inc
          PO Box 3633
          Bellevue, WA 98009-3633, USA
          Phone:  +1 (206) 827-3304 (800-676-7568 in US/Canada)

          Fax:    +1 (206) 827-3989
          E-Mail: tecplot@amtec.com
          Contact: Tom Chan

Entered:

Sep 24, 1994

5.3 Development tools

Compilers, development environments and so on.

5.3.1 Basmark QuickBASIC

Description:

The Basmark QuickBASIC Compiler is a multi-user IBM-PC BASICA, MBASIC and Microsoft Quick-BASIC Compiler designed to provide performance and consistency across a variety of machines (e.g. i386 and i486, Pentium, SPARC, RS/6000, HP PA-Risc) under Unix, AIX, SunOS, Linux, HP-UX, Xenix.

Distribution Media:

3.5” diskette Licensing Policy:

Per machine, no run-time restrictions. OS Provisions:

GNU GAS and LD must be installed. Documentation:

400+ page manual, release and installation notes. Product Support:

Updates available for $39.00 less shipping. Maintenance contracts available. Linux Support:

N/A Extra Features and Add-Ons:

C-ISAM (Informix Inc.) Interface in C source code form. Cost $35.00. Available Since:

December 1993 Countries with Distribution:

to purchase from manufacturer directly.

Price Range:

$195.00 (less shipping) Installed Base: Vendor:

     Basmark Corporation
     P.O. Box 40450
     Cleveland, OH 44140, USA
     Phone:  +1 (216) 871-8855
     Fax:    +1 (216) 871-9011
     E-Mail: jgo@ios.com  (for orders)
     Contact: Joseph O'Toole (for orders)

Entered:

Sep 23, 1994

5.3.2 ISE Eiffel 3

Description:

ISE Eiffel 3 provides a powerful and user-friendly O-O programming environment designed for large, complex systems. It is an integrated GUI workbench consisting of a variety of Eiffel-based components: EiffelBench melting-ice workbench, EiffelBuild interface builder and application generator, EiffelVision graphics and GUI library, and EiffelBase basic libraries.

Distribution Media:

1/4 inch Tape, DAT. Licensing Policy:

Each hardware platform type requires a different runtime. The standard delivery of the EiffelBench comes with the runtime for the given platform type; additional runtimes may be purchased for crossdevelopment. ISE charges run-time license fees, only for copyrighted commercial products built using ISE’s technology. For further information contact vendor.

OS Provisions:

Motif required Documentation:

Documentation including the Prentice Hall books "Eiffel: The Language, Reusable Software: The Base Object-Oriented Component Libraries" and "An Object-Oriented Environment: Principles & Practices".

Product Support:

The purchase prices include free technical support for three months. The price of technical support per year is 30% (payable in advance) of the standard software purchase and provides assistance with problems encountered in the use of the software. Support is provided by electronic mail, fax or telephone.

Linux Support:

N/A

Extra Features and Add-Ons:

A special license for students and individuals of ISE Eiffel on Linux includes the full graphical environment.

Available Since:

June 1994

Countries with Distribution:

(many individual distributors worldwide; too many to name them here. please consult vendor. hm)

Price Range:

US$295.00 plus shipping & handling. USA: $15.00 for UPS Ground, $20.00 for UPS 2nd Day Air; Canada: $35.00 for UPS 2nd Day Air; Other: $60.00 for UPS International Air

Installed Base:

Vendor:

     Interactive Software Engineering, Inc.
     270 Storke Road, Suite 7
     Goleta, CA 93117, USA
     Phone:  +1 (805) 685-1006
     Fax:    +1 (805) 685-6869
     E-mail: queries@eiffel.com

Entered:

Sep 09, 1994

5.3.3 Metacard

Description:

MetaCard is a hypermedia/Rapid Application Development environment for X11/Unix workstations that is compatible with Apple Corp.’s HyperCard. MetaCard can be used by programmers and sophisticated end users to build Motif applications and hypermedia documents using a powerful, direct manipulation editor and a simple scripting language. Stacks developed with MetaCard are portable among all supported platforms (14 for release 1.3) and can be distributed with the MetaCard engine without licensing fees or royalties.

Distribution Media:

Anonymous FTP and 3.5” diskettes. Licensing Policy:

MetaCard can be licensed to a single, named individual, but can be used on any machine or combination of machines by that user. Multiple user packages are also available. The save-disabled distribution is available via anonymous FTP from ftp.metacard.com (ftp://ftp.metacard.com/MetaCard) and

ftp.uu.net (ftp://ftp.uu.net/vendor/MetaCard).

OS Provisions:

The 1.4 release was built on Linux 0.99pl14. 8 MByte RAM minimum, 16MByte recommended. 800x600 minimum display resolution. Requires about 5 MByte disk space.

Documentation:

Complete on-line documentation is supplied. Printed documentation is an extra-cost option. Product Support:

Free email technical support. Free upgrades during the 1.X series of releases. Linux Support:

N/A Extra Features and Add-Ons:

A library version of MetaCard that can be linked directly to C programs (Embedded MetaCard) is available at extra cost.

Available Since:

February 1994 Countries with Distribution:

Direct sales world wide. Price Range:

$247.50. $495 for a license that works on Linux and all Unix platforms. Installed Base: Vendor:

     MetaCard Corporation
     4710 Shoup pl.
     Boulder, CO 80303, USA
     Phone:  +1 (303) 447-3936
     Fax:    +1 (303) 499-9855
     E-Mail: info@metacard.com
     Contact: Scott Raney

Entered:

Sep 20, 1994

5.3.4 tgdb Description:

tgdb is a graphical user interface for gdb, the GNU debugger. Distribution Media:

Anonymous  ftp,  available

     as source (ftp://sunsite.unc.edu:/pub/Linux/devel/debuggers/tgdb-1.0.src.tgz) package or

Linux binary

     (ftp://sunsite.unc.edu:/pub/Linux/devel/debuggers/tgdb-1.0.srcbin-linux.tgz) package.

Licensing Policy:

Shareware OS Provisions:

Any Un*x-like OS (including Linux); requires X11R5. Documentation:

Comprehensive online/hypertext help (english); can be converted into various formats for printing. Product Support:

Updated when required; hotline and email support available. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

August 1994 Countries with Distribution:

worldwide (ftp) Price Range:

US$30 or DM50.00 per user; site licenses available. Installed Base: Vendor:

     HighTec EDV-Systeme GmbH
     Neue Bahnhofstr. 37
     D-66386 St. Ingbert, Germany
     Phone: +49 (0)6894/87 00 41
     Fax:   +49 (0)6894/87 00 44
     Email: tgdb@hightec.saarlink.de
     Contact: Michael Schumacher

Entered:

Sep 09, 1994

5.4 Text Processing

All kind of text processing, ASCII- and GUI-based

5.4.1 CRISP

Description:

CRISP is a graphical text editor on various Unix and Windows platforms which is 100% compatible with BRIEF. What makes Crisp different is that it tries to deliver its power in an intuitive point & click environment, without taking away the keyboard.

Distribution Media:

3.5” diskettes. Tapes available at extra charge. Licensing Policy:

Both node locked and floating licenses are available. Demonstration copy is archived at ftp.uu.net

(ftp://ftp.uu.net/vendor/vital).

OS Provisions:

RAM: 0.5 MByte, harddisk: 6 MByte (full), 2 MByte minimum. Openlook / Motif / Character versions are available on just about all commercial Unix platforms, Linux and Windows 3.1/NT.

Documentation:

printed, on-line manuals, on-line help, English Product Support:

updates and maintenance can be purchased at $100.00 US per license per year. It provides free upgrades for one year.

Linux Support:

We are currently giving away Slackware Professional 2.0 away for users ordering Crisp. This promotion however is only valid till end of October.

Extra Features and Add-Ons:

N/A Available Since:

March 1994 Countries with Distribution:

International Distributor: VITAL; Europe: Lasermoon Price Range:

$99.99 without printed manuals. Printed user manual can be purchased at an additional $40.00 Installed Base: Vendor:

          VITAL SOLUTIONS INC
          4109 Candlewyck Drive
          Plano, TX 75024, USA
          Phone:  +1 (214) 491-6907
          Fax:    +1 (214) 491-6909
          E-Mail: info@vital.com
          Contact: Gigi Mehrotra

Entered:

Sep 20, 1994

5.4.2 ibgsXaed

Description:

ibgsXaed is an editor which is portable to every X Window System that offers OSF/Motif. The editor has normal edit functions and a lot of special features that support newcomers and professionals (f.e. function keys for system commands execution, variable fonts, undo, redo, online help, hotlist, iconbar, etc.).

Distribution Media:

3.5” diskette Licensing Policy:

free demo version OS Provisions:

X Window System X11R5, OSF/Motif 1.2, approx. 2 MByte harddisk Documentation:

printed manual, online help, both in English or German Product Support:

hotline, updates Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1994 Countries with Distribution: Price Range:

DM 799,- for one license, sliding prices for more licenses Installed Base: Vendor:

          ibgs GmbH
          Promenade 7
          D-52076 Aachen
          Phone:  +49 (0)2408/9455-53
          Fax:    +49 (0)2408/9455-15
          Contact: Mr.Sohn

Entered:

Sep 28,1994

5.5 X Windows related stuff

Alternative X Windows servers, widget sets, window managers.

5.5.1 Metrolink Motif

Description:

Motif 1.2.4 - Graphical User Interface - Complete Runtime and Development package with Color Pixmap Support.

Distribution Media:

3.5” diskettes (3), or FTP: The customer places the order with a credit card, then we email him our license agreement, installation instructions, release notes, and the FTP instructions – including the current password.

Licensing Policy:

Licensing is per CPU. NFS mounting not allowed. Site licenses are available for 10 or more copies. OS Provisions:

Linux 0.99pl15g or higher, 8-16 MByte RAM, 9 MByte HardDisk Documentation:

OSF User’s Manual (120 pages), man pages (electronic), English. Product Support:

90 days free tech support, 30-day money back guarantee, Extended Tech Support Contract $100 US per year.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

Motif 1.2.4 for Linux since August 1994 (Motif 1.2.2 for Linux since June 1993). Countries with Distribution:

USA, Germany, France. Price Range:

$149.00 US through October 31, 1994; $199 US after that if the special offer isn’t extended. Installed Base:

2,600 copies as of Sep 23, 1994 Vendor:

          Metro Link Incorporated
          4711 N. Powerline Rd.
          Fort Lauderdale, FL 33309, USA
          Phone:  +1 (305) 938-0283
          Fax:    +1 (305) 938-1982
          E-Mail: sales@metrolink.com (general info)
          Contact:  Holly Robinson (holly@metrolink.com, other questions \& orders)

Entered:

Sep 19, 1994 5.5.2 X Inside X Servers

Description:

Replacement X Server for the XFree86 and other Servers for Linux. Supports Actix, ColorGraphics, Compaq, ELSA, Matrox, Number 9, ATI, Boca, Diamond, Orchid, Reveal, STB, TechWorks. Chipsets include Matrox, Number 9 Imagine-128, ATI Mach64, S3 964. Supports Mouse Systems, Microsoft, Logitech mice.

Distribution Media:

single floppy diskette Licensing Policy:

Per cpu license. Site licenses available. OS Provisions:

Tested on versions of Linux between 1.0.8 and 1.1.18. No known problems with any Linux distribution. Requires around 3MB hard disk. Best used with a mouse...

Documentation:

Postscript formatted manual available via ftp (ftp://ftp.xinside.com/accelx/1.0/Manual). Hardcopy shipped with first copy of Server license.

Product Support:

Installation support included in price; available by email, Fax (and by phone if we have to). Update service available, US$100 for four revisions a year, to track new chipsets and other performance improvements. US$50 to upgrade to specified revision (e.g. 1.1 to 1.2). Product automatically sends email on failure. We may respond to a problem before you know you have it!

Linux Support:

N/A Extra Features and Add-Ons:

Various extensions will be offered, starting with PEX. Available Since:

May 94 Countries with Distribution:

Distribution currently in negotiation. Complete information available via ftp (

ftp://ftp.xinside.com/accelx/1.0/All/prodinfo.txt).

Price Range:

US$199.00. Student, academic and VAR discount programs. Student discount is 20% for Release 1.0 and 50% for Release 1.1. Post & Packing is between US$3 for US Mail to US$15 for Next Day delivery.

Installed Base:

Vendor:

     X Inside Incorporated
     P.O. Box 10774
     Golden, CO 80401-0610, USA
     Phone:  +1 (303) 384-9999
     Fax:    +1 (303) 384-9778
     E-mail: info@xinside.com

Entered:

Sep 09, 1994

5.6 Other Software

Everything else that does not fit in one of the other sections by now.

5.6.1 aqua zis

Description:

(Time Series Information System). A system to store and maintain huge amounts of measurement data (mainly hydrological data). Time series with several 100,000’s of values can be handled. You may retrieve, edit and insert data very fast. A lot of statistical and arithmetical operators are applicable. Time series may be visualized in a full scale GUI, includeing high performance zooming ans scrolling. The user may define individual GUI’s in an easy script language and with the time series programming language Azur. Reports on any paper size including full graphics are possible through Azur. Output is produced in Postscript and HPGL formats.

Distribution Media:

We install at your site. Demo disk available (tar)

Licensing Policy:

machine or site licenses available. Test licenses are limited in time.

OS Provisions:

Linux version 0.99pl6 or later. Motif required if you use the GUI features. When set up as a time series database server, no graphics is needed. RAM usage 8 – 16 MB. Disk space required: some 15 MB for executables and help files. For the data: several GB’s.

Documentation:

printed manuals covering end user and admin tools (English version soon to appear). Program documentation and design available only in English.

Product Support:

12 months warranty, maintenance contract recommended. One day training at your site included.

Linux Support:

N/A Extra Features and Add-Ons:

aqua trop precipitaion reconstrucion tool using GIS methods. DigiTul digitizing of precipitaion, waterlevel etc. paper registrations. aqua log gound water data management.

Available Since:

Jan 1994 Countries with Distribution:

Germany, Austria, soon EC. Price Range:

machine license: 39,000 DM, multiple machines: graduated discounts, site licenses on request, reasonable discout for universities.

Installed Base:

some 4 Linux-only systems are running. Vendor:

     aqua_plan
     Ing.-Ges. fuer Problemloesungen
     in Hydrologie und Umweltschutz mbH
     Mozartstr. 16
     D-52064 Aachen, Germany
     Phone:  +49 (0)241/31430
     Fax:    +49 (0)241/31499

Entered:

Aug 26, 1994

5.6.2 Maple V

Description:

Maple V Release 3 is a powerful general purpose computer algebra system. Maple V is a system for solving mathematical problems symbolically (instead of using paper and pencil and a lot of time) and numerically. Maple V’s graphics (various types of 2D, 3D, animation) visualizes the solutions. Maple V has also a programming language (Pascal like) which allows to extend the library of 2500+ functions.

Distribution Media:

3.5” disks Licensing Policy:

License for unlimited use in time of the bought version, floating license with user based pricing OS Provisions:

Linux kernel 1.0; XFree86 or Motif possible but not necessary; release number 2100: at least 8 MByte RAM suggested

Documentation:

3 books (”First Leaves: A Tutorial Introduction” 253 pgs., ”Language Reference Manual”, 267 pgs. and ”The Maple Handbook” 497 pgs.), Installation Manual. Manual pages of all functions and commands, help browser, keyword search. Documentation and help in English, some books in various languages exist.

Product Support:

updates/maintenance depending on license, please call the distributor or WMSI.

Linux Support:

N/A

Extra Features and Add-Ons:

share library via email and ftp (ftp://daisy.uwaterloo.ca/pub/maple). Free mailing list available

(maple group@daisy.waterloo.edu).

Available Since:

September 94

Countries with Distribution:

US/Canada, Europe, Asia

Price Range:

German list price: DM 1785 (plus shipping and handling plus VAT). Please call the national distributor or WMSI.

Installed Base:

Vendor:

     Waterloo Maple Software
     450 Phillip Street, Waterloo, Ontario, Canada N2L 5J2
     Phone: +1 (519) 747-2373
     Fax:   +1 (519) 747-5284
     email: info@maplesoft.on.ca
     Germany, Switzerland and Austria:
     Scientific Computers GmbH
     Franzstr. 106
     D-52064 Aachen, Germany
     Phone: +49 (0)241/26041
     Fax:   +49 (0)241/44983
     E-Mail: info@scientific.de
     Contact: Monika Germ

Entered:

Sep 19, 1994

5.6.3 Reduce

Description:

REDUCE is a general purpose system for the symbolic manipulation of mathematical formulae (computer algebra) in science and engineering. More information via WWW

(http://www.zib-berlin.de/Symbolik/reduce).

Distribution Media:

REDUCE for Linux is normally distributed on 3.5 inch diskettes. Tape cartridges of different formats are available on request.

Licensing Policy:

One time fee per installation. Discount for multiple licenses. Free demo version available via ftp

(ftp://ftp.zib-berlin.de/pub/reduce/demo/linux).

OS Provisions:

Independent of LINUX kernel: the .c and .s sources for the LISP kernel are included and can be recompiled locally. X Windows not necessary but recommended. 8 Mbyte RAM recommended, better more.

Documentation:

A 230 page User’s manual is available. A reference manual is accessible in hypertext style on line under X Windows. All documents in English. The manual and additional documentation on disk in LaTeX.

Product Support:

Free patches of the current version available via electronic mail or ftp. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1993 Countries with Distribution:

All countries except Iraq and Yugoslavia. Local distributors in Russia, Japan, Czec Republic, China. Price Range:

Personal REDUCE: approx. US$ 99; Professional REDUCE: approx. US$ 499. Plus freight and tax.

Installed Base:

several dozens

Vendor:

     Konrad-Zuse-Zentrum Berlin
     Heilbronner Str. 10
     D-10711 Berlin-Wilmersdorf, Germany
     Phone:  +49 (0)30/89604-195
     Fax:    +49 (0)30/89604-125
     E-Mail: melenk@zib-berlin.de
     Contact: H.Melenk

Entered:

Sep 28, 1994

5.6.4 ROSIN NC

Description:

ROSIN-NC-Werkbank (ger: workbench) The ROSIN-NC-Werkbank is a comprehensive set of tools which accomplish a CAD system. These tools comprise the ROSIN-VDAFS-Processor to process digitised data, and data in the VDAFS format; the ROSIN-Post Processors to generate 2- to 5-axis milling programs for all major CNC process controls; the ROSIN-NC-Editor to analyse and edit existing NC programs; the ROSIN-NC-Visualiser to visualise the tool path and generate printed information about a NC program; the ROSIN-Engraver to mill logos or fonts of any kind.

Distribution Media:

Tapes (DAT, QuarterInch) or floppy disk

Licensing Policy:

All tools of the ROSIN-NC-Werkbank are independently licensed via the ROSIN License System, the customer gets a production license for each purchased tool. The combination of license types with different tools is possible. There are individual licenses for single users, 2-5 users etc. Please consult vendor for more information.

OS Provisions:

Kernel Version higher or equal to 0.99pl14; Motif; RAM 16 MByte, recommended 20 MByte; the necessary disk space for a complete installation with all tools is 25 MByte. It is recommended to ensure a free disk space of at least 1,5 times the size of the input file(s).

Documentation:

An Online Help Function is available with all tools. With the purchase of each tool a manual is supplied in the language chosen by the customer (languages available: German, English, French and Italian, other languages on request). Size of manuals: Depending on the tool from 60 to 170 pages each.

Product Support:

Maintenance contracts are offered for all tools and cost 1,3% per month of the license price. With a maintenance contract the customer gets regular product updates and manual updates. Migration to a different hardware platform is significantly cheaper (3% of the license price instead of 12% for customers without maintenance contract). The telephone hotline is free of charge. Our telephone hotline is available from 8.30 to 17.00 MET, languages are German and English. We offer special prices for product installation and training courses (75% of the normal service prices) for product customers. Prices for product updates can be provided on request.

Linux Support:

(there was no ”official” statement by ROSIN, however, I was told on the phone that ROSIN ships complete systems with pre-installed software, and that they would offer limited Linux support as far as handling their software is concerned. hm)

Extra Features and Add-Ons:

N/A

Available Since:

1994

Countries with Distribution:

Germany, France and Mexico (please consult vendor for more information). We are looking for distributors in the UK, Spain and Italy.

Price Range:

The lower prices (in DM) are for Single User Licenses, the higher ones for Multi User Licenses, floating type. Prices for all other licenses are included in this price range. ROSIN-VDAFS-Processor 7500 to 9750 ROSIN-3-axis Post Processor 4500 to 11050 ROSIN-5-axis Post Processor 13500 to 32500 ROSIN-NC-Editor 7500 to 9750 ROSIN-NC-Visualiser 3500 to 4550 ROSIN-Engraver 1500 to 1950

Installed Base:

Vendor:

     ROSIN Gesellschaft fuer technische Datenverarbeitung mbH
     Roniger Weg 13
     53545 Linz am Rhein, Germany
     Phone:  +49 (0)2644/97003-0
     Fax:    +49 (0)2644/97003-32
     Contact: Gisela Buechner (Marketing)

5.7 Free Software for Commercial Hardware

This section is somewhat ”exotic”. It covers free Linux software (drivers et cetera) for specialized hardware. This is no ”commercial software”, however, the software is unusable without the hardware.

5.7.1 Cyclades Cyclom-Y

Description:

Cyclom-Y is serial multiport card for ISA/EISA PC’s. 8 and 16-port models are available. There are RJ-12 and DB-25 connector options. It is based on the Cirrus Logic CD1400 serial processor. This RISC-based processor supports on-board character processing and allows speeds up to 115.2 kbps. Drivers for Xenix, SCO, ISC and UnixWare are available from Cyclades. There is a driver for Linux available in the Internet. Board variants: Cyclom-8Ys: 8 ports, RJ-12 connectors (no RTS signal); Cyclom-8Yo: 8 ports, DB-25 connectors in a octopus cable (full-modem); Cyclom-8Yb: 8 ports, DB-25 connectors in a external box (full-modem); Cyclom-8Yb+: 8 ports, same as 8Yb with surge protection; Cyclom-16Y: 16 ports, DB-25 connectors in a external box (full-modem).

Distribution Media:

3.5” diskettes

Licensing Policy:

The Linux device driver software is free and is available via ftp at sunsite.unc.edu

(ftp://sunsite.unc.edu/pub/Linux/kernel/patches/misc).

OS Provisions:

(the driver is available as a patch relative to Linux 1.1.8. A driver for newer kernels is under development. hm)

Documentation:

The user’s manual contains all the reference information to configure and install the card. Specific information about the Linux driver can be found in readme files.

Product Support:

SW updates from Internet. Cyclades will provide technical support on the hardware, but the driver is a freeware and is provided ”as is”. Cyclades makes no warranties about the suitability of this software for any purposes. Some companies bundle the driver in their Linux distributions and can help in the installation and configuration of the card.

Linux Support:

N/A

Extra Features and Add-Ons:

N/A

Available Since:

July 1994 5. Product Groups 35

Countries with Distribution:

worldwide

Price Range:

List Price (end user):
CYCLOM 8YS \$459
CYCLOM-8YO \$549
CYCLOM-8YB \$599
CYCLOM-8YB+ \$699
CYCLOM-16Y/DB25 \$794
Promotion for resellers only (first-time buyer):
CYCLOM 8YS \$99
CYCLOM-8YO \$199
CYCLOM-16Y/DB25 \$399
NOTE:  All prices are F.O.B. Fremont, CA, USA

Installed Base:

Vendor:

Cyclades Corporation.
44140 Old Warm Springs Blvd
Fremont, CA 94538, USA
Phone:  +1 (510) 770-9727 or (800) 347-6601
Fax:    +1 (510) 770-0355
E-Mail: cyclades@netcom.com
Contact: Doris Li (Sales and Marketing).

Entered:

Sep 24, 1994